# MUCAR: Benchmarking Multilingual Cross-Modal Ambiguity Resolution for Multimodal Large Language Models

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant advances across numerous vision-language tasks. Due to their strong image-text alignment capability, MLLMs can effectively understand image-text pairs with clear meanings. However, effectively resolving the inherent ambiguities in natural language and visual contexts remains challenging. Existing multimodal benchmarks typically overlook linguistic and visual ambiguities, relying mainly on unimodal context for disambiguation and thus failing to exploit the mutual clarification potential between modalities. To bridge this gap, we introduce MUCAR, a novel and challenging benchmark designed explicitly for evaluating multimodal ambiguity resolution across multilingual and cross-modal scenarios. MUCAR includes: (1) a multilingual dataset where ambiguous textual expressions are uniquely resolved by corresponding visual contexts, and (2) a dual-ambiguity dataset that systematically pairs ambiguous images with ambiguous textual contexts, with each combination carefully constructed to yield a single, clear interpretation through mutual disambiguation. Extensive evaluations involving 19 state-of-the-art multimodal models—encompassing both open-source and proprietary architectures—reveal substantial gaps compared to human-level performance, highlighting the need for future research into more sophisticated cross-modal ambiguity comprehension methods, further pushing the boundaries of multimodal reasoning.
@@ 1 Introduction
To systematically evaluate the capabilities of MLLMs to resolve such complex multimodal ambiguities, we introduce MUCAR, a novel benchmark specifically designed for MUltilingual Cross-modal Ambiguity Resolution. Table 1 summarizes recent benchmarks designed to evaluate MLLMs in terms of visual ambiguity, contextual ambiguity, multilinguality, and the type of evaluator used (e.g., metrics, GPT, or human annotations). While early benchmarks such as MMT-Bench (Ying et al., 2024), MMStar (Liu et al., 2023c), and MME-RealWorld (Li et al., 2023) focus on general multimodal tasks, they lack coverage of ambiguity-related phenomena. More recent benchmarks like HallusionBench (Guan et al., 2023), Illusory VQA (Rostamkhani et al., 2024), and CODIS (Luo et al., 2024) begin to explore visual ambiguity, but often overlook contextual disambiguation or multilingual diversity. Notably, only a few benchmarks incorporate human evaluation, which is essential for assessing ambiguity understanding. To the best of our knowledge, MUCAR is the first benchmark to comprehensively address visual ambiguity, contextual ambiguity, and multilinguality, while integrating both human and GPT-based evaluation. This design enables a more rigorous and realistic assessment of ambiguity resolution capabilities in multimodal large language models.
@@ 3.1 Taxonomy
Polysemy. Navigli (2009) defines the term, polysemy, by referring to a word with two or more related meanings. These related meanings often share a conceptual core, with one meaning typically being an extension or variation of the other. Some contexts can be interpreted both literally and metaphorically. The “Polysemy” part in Figure 2 gives a good example.
Homonymy. Homonymy is also defined in Navigli (2009). Opposite to polysemy, homonymy refers to a word having two or more unrelated meanings that stem from different historical origins, and the meanings of homonyms have no inherent connection. Disambiguating homonyms relies on the other elements. The “Homonymy” part in Figure 2 serves as a good example.
{CJK*}
UTF8gbsn
@@
Specialized. Context can be interpreted differently depending on the circumstances of the viewer, including their background knowledge, expertise, or personal situation. What is clear and unambiguous to someone with specialized training might be confusing or misleading to a novice. The “Specialized” part in Figure 2 presents a good example.
@@ 3.2 Instruction Design
MLLMs receive $i$ queries independently and produce their outputs $(\mathcal{O}_{1},\mathcal{O}_{2},\dots,\mathcal{O}_{i})$. These outputs are evaluated by comparing them with ground truth answers $(\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{i})$.
@@ 3.4 Evaluation Method
Open-source Models $>7\mathrm{B}$
@@
Open-source Models $>7\mathrm{B}$
@@
Table 2: Results of MLLMs on MUCAR benchmark under two settings: (top) without confusing options and (bottom) with confusing options.
@@
where $n_{k}$ denotes the number of queries within each pair, $n_{p}$ represents the total number of data pairs, and $n_{q}$ is the total number of individual queries. For $\mathrm{Acc_{p}}$, a model receives a score only if both answers within a pair are correct. For $\mathrm{Acc_{q}}$, a model scores for each correctly answered query.
@@ 4.1 Models
We evaluate a total of 19 models covering a range of scales and architectures. Our evaluated proprietary models include GPT-4V (OpenAI, 2023), GPT-4o (OpenAI, 2024), Gemini (Gemini Team et al., 2023), and Claude-3.5-Sonnet (Anthropic, 2024). For open-source models, we include Deepseek-VL-Tiny (Lu et al., 2024), Kimi-VL (Team et al., 2025), Llama-3.2-Vision-11B (Meta, 2024), MiniCPM-o 2.6 (Yao et al., 2024), InternVL2.5 series (Wang et al., 2024b; Chen et al., 2024c), LLaVA-v1.6-vicuna-7b (Liu et al., 2024), Qwen2.5-VL series (Team, 2025). Details of these models are listed in Table 7 in Appendix B.
@@ 4.2 Main Results
Overall Performance. Across all evaluated models, InternVL2.5-8B-MPO-AWQ achieves the best overall accuracy (Accq = 67.03%), followed closely by MiniCPM-o 2.6 (66.48%) and InternVL2.5-8B-MPO (65.71%). Among proprietary models, GPT-4o outperforms the others, obtaining an overall accuracy of 53.91%, slightly higher than GPT-4V (52.70%). In contrast, Claude-3.5-Sonnet and Kimi-VL underperform, showing limited ability in disambiguation tasks.
Results with Different Model Size. Open-source models with scales larger than 7B generally outperform smaller ones, with all top-performing models falling within this range of scale, which is likely to benefit from richer training data and more advancedadvanced architectures. In comparison, models smaller than or equal to 7B show a clear performance gap. Although certain models, such as LaVA-Onevision (61.30%) and mPLUG-Owl2-3B (58.21%), perform competitively, most smaller models struggle with complex ambiguities, particularly in semantic and cultural contexts.
Results on Different Categories. We further break down the results by disambiguation categories, and find that InternVL2.5-8B-MPO-AWQ consistently leads in most categories, especially in Homonymy (70.18%), Grammar (64.71%), and Semantics (61.90%). Notably, MiniCPM-o 2.6 excels in the Specialized category (69.47%), suggesting domain knowledge plays a key role. In the Cultural category, which requires understanding cross-cultural references, models like GPT-4o (34.12%) and InternVL2.5-8B-MPO-AWQ (60.00%) show relatively stronger performance. On the other hand, most models perform poorly in the Polysemy and Dual-Ambiguity categories, reflecting the inherent challenges in resolving subtle or cross-modal ambiguities.
@@
5 Analyses and Discussion
5.1 Ablation on Benchmark
@@ 5.2 Discussion
5.3 Further Exploration: An Agent Framework for Ambiguity Resolution
Figure 5: Pipeline of our proposed framework.
@@ 5.3 Further Exploration: An Agent Framework for Ambiguity Resolution
Formally, given a task description $\mathcal{D}$, a question $\mathcal{Q}$, and a multimodal ambiguous context $\mathcal{X}$ (e.g., an image and textual), the agent-based model $\mathcal{M}$ solves the task through a three-step process as illustrated in Figure 5:
@@ 6 Conclusion
Appendix A Details of Datasets
CODIS (Luo et al., 2024) A benchmark to assess the capabilities of MLLMs in context-dependent visual comprehension, consisting of 706 queries over 216 images across five categories.
MMA (Wang et al., 2024a) A multiple-choice VQA benchmark focused on ambiguity, including 261 questions paired with two divergent images that yield different answers given the same textual context.
@@ Appendix A Details of Datasets
Table 6 presents the detailed prompts used during model testing. In the main experiments, we employed English prompts; additionally, we conducted ablation studies using Chinese and Malay prompts to evaluate the impact of different evaluation languages on the experimental results. The table also separately shows the prompts used in the three ablation settings: (Q, I), (C, Q), and C. {CJK*}UTF8gbsn
@@
Appendix B Evaluated Models
We evaluate a total of 19 models covering a range of scales and architectures. Our evaluated proprietary models include GPT-4V (OpenAI, 2023), GPT-4o (OpenAI, 2024), Gemini (Gemini Team et al., 2023), and Claude-3.5-Sonnet (Anthropic, 2024). For open-source models, we include Deepseek-VL-Tiny (Lu et al., 2024), Kimi-VL (Team et al., 2025), Llama-3.2-Vision-11B (Meta, 2024), MiniCPM-o 2.6 (Yao et al., 2024), InternVL2-8B (Chen et al., 2024d), InternVL2.5-8B-MPO (Wang et al., 2024b), InternVL2.5-8B-MPO-AWQ (Chen et al., 2024c), LLaVA-v1.6-vicuna-7b (Liu et al., 2024), Qwen2.5-VL-3B-Instruct (Team, 2025), Qwen2.5-VL-7B-Instruct (Team, 2025). Details of these models are listed in Table 7. Table 7 presents a comprehensive overview of the Multimodal Large Language Models (MLLMs) evaluated in our benchmark. The models are categorized into two groups: API-based models and open-source models. For each model, we list its parameter size category (greater than or less than 7 billion), the vision encoder architecture, the underlying language model (LLM) backbone, and the employed vision-to-language (V2L) adapter. API-based models such as GPT-4V and Gemini do not publicly disclose architectural details, while open-source models span a variety of encoders (e.g., SigLIP, CLIP ViT, InternViT), LLM backbones (e.g., Llama, Qwen, InternLM), and adapter types (e.g., MLP, Linear, XAttn LLM). This table highlights the diversity in architectural design choices across MLLMs.
@@ Appendix B Evaluated Models
Deepseek-VL-Tiny (Lu et al., 2024) $>7\mathrm{B}$ SigLIP DeepSeek LLM MLP
@@
Qwen2.5-VL-3B-Instruct (Team, 2025) ViT Qwen2.5 LLM MLP
Qwen2.5-VL-7B-Instruct (Team, 2025) ViT Qwen2.5 LLM MLP
@@
Table 8 presents the detailed experimental results of multiple models with the language set to Chinese, conducted without any interference options enabled.
@@
Table 8: We input prompts in Chinese, and the options are without confusing elements.
Table 9 shows the detailed experimental results of multiple models evaluated with the language set to Chinese, conducted with interference options enabled.
Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq
@@
Table 9: We input prompts in Chinese, and the options are with confusing elements.
Table 10 provides the detailed experimental results of multiple models evaluated with the language set to Malay, conducted without any interference options enabled.
Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq
@@
Table 10: We input prompts in Malay, and the options are without confusing elements.
Table 11 presents the detailed experimental results of multiple models evaluated with the language set to Malay, conducted with interference options enabled.
Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq Accp Accq
@@
Table 11: We input prompts in Malay, and the options are with confusing elements.
Appendix C More Cases

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated significant advances across numerous vision-language tasks. MLLMs have shown promising capability in aligning visual and textual modalities, allowing them to process image–text pairs with clear and explicit meanings. However, resolving the inherent ambiguities present in real-world language and visual contexts remains a challenge. Existing multimodal benchmarks typically overlook linguistic and visual ambiguities, relying mainly on unimodal context for disambiguation and thus failing to exploit the mutual clarification potential between modalities. To bridge this gap, we introduce MUCAR, a novel and challenging benchmark designed explicitly for evaluating multimodal ambiguity resolution across multilingual and cross-modal scenarios. MUCAR includes: (1) a multilingual dataset where ambiguous textual expressions are uniquely resolved by corresponding visual contexts, and (2) a dual-ambiguity dataset that systematically pairs ambiguous images with ambiguous textual contexts, with each combination carefully constructed to yield a single, clear interpretation through mutual disambiguation. Extensive evaluations involving 19 state-of-the-art multimodal models–encompassing both open-source and proprietary architectures–reveal substantial gaps compared to human-level performance, highlighting the need for future research into more sophisticated cross-modal ambiguity comprehension methods, further pushing the boundaries of multimodal reasoning.
@@ 1 Introduction
To systematically evaluate the capabilities of MLLMs to resolve such complex multimodal ambiguities, we introduce MUCAR, a novel benchmark specifically designed for MUltilingual Cross-modal Ambiguity Resolution. Table 1 summarizes recent benchmarks designed to evaluate MLLMs in terms of visual ambiguity, contextual ambiguity, multilinguality, and the type of evaluator used (e.g., metrics, GPT, or human annotations). While early benchmarks such as MMT-Bench (Ying et al., 2024), MMStar (Liu et al., 2023c), and MME-RealWorld (Li et al., 2023) focus on general multimodal tasks, they lack coverage of ambiguity-related phenomena. More recent benchmarks like HallusionBench (Guan et al., 2023), Illusory VQA (Rostamkhani et al., 2024), and CODIS (Luo et al., 2024) begin to explore visual ambiguity, but often overlook contextual disambiguation or multilingual diversity. Notably, only a few benchmarks incorporate human evaluation, which is essential for assessing ambiguity understanding. To the best of our knowledge, MUCAR is the first benchmark to comprehensively address visual ambiguity, contextual ambiguity, and multilinguality, while integrating both human and GPT-based evaluation. This design enables a more rigorous and realistic assessment of ambiguity resolution capabilities in multimodal large language models. Our dataset and code are available at https://github.com/THUNLP-MT/MUCAR.
@@ 3.1 Taxonomy
Polysemy. Navigli (2009) provides one formalization of polysemy, by referring to a word with two or more related meanings. These related meanings often share a conceptual core, with one meaning typically being an extension or variation of the other. Some contexts can be interpreted both literally and metaphorically. The “Polysemy” part in Figure 2 gives a good example.
Homonymy. Navigli (2009) also provides one formalization of homonymy. Opposite to polysemy, homonymy refers to a word having two or more unrelated meanings that stem from different historical origins, and the meanings of homonyms have no inherent connection. Disambiguating homonyms relies on the other elements. The “Homonymy” part in Figure 2 serves as a good example.
@@
Specialized. We define Specialized taxonomy to encompass terms or concepts that have distinct meanings across academic domains or personal situations. These terms often lead to ambiguity when encountered by individuals with varying background knowledge. The “Specialized” part in Figure 2 presents a good example.
@@ 3.2 Instruction Design
Each $(\mathcal{C},\mathcal{Q},\mathcal{I})$ is independently input into MLLMs as a query without being influenced by other queries in the same pair. MLLMs receive $i$ queries in the same pair independently and produce their outputs $(\mathcal{O}_{1},\mathcal{O}_{2},\dots,\mathcal{O}_{i})$. These outputs are evaluated by comparing them with ground truth answers $(\mathcal{A}_{1},\mathcal{A}_{2},\dots,\mathcal{A}_{i})$.
@@ 3.4 Evaluation Metrics
Open-source Models $7\mathrm{B}$
@@
Open-source Models $7\mathrm{B}$
@@
Human
Human 75.00 86.59 73.50 86.00 70.09 85.05 78.57 89.29 79.55 89.36 68.29 83.53 81.43 87.12 74.66 86.42
Table 2: Results of MLLMs on MUCAR benchmark under two settings: (top) without confusing options and (bottom) with confusing options. For humans, the two experimental settings show little difference, so we chose the more challenging with confusing options setting for human evaluation.
@@
where $n_{k}$ denotes the number of queries within each pair, $n_{p}$ represents the total number of pairs, and $n_{q}$ is the total number of individual queries. $\mathrm{Acc_{p}}$ denotes the accuracy of judging each individual query’s correctness independently. $\mathrm{Acc_{q}}$ requires that a pair is considered correct only if MLLMs correctly judge all queries within the pair.
@@ 4.1 Models
We evaluate a total of 19 models covering a range of scales and architectures. Our evaluated proprietary models include GPT-4V (OpenAI, 2023), GPT-4o (OpenAI, 2024), Gemini (Gemini Team et al., 2023), and Claude-3.5-Sonnet (Anthropic, 2024). For open-source models, we include Deepseek-VL-Tiny (Lu et al., 2024), Kimi-VL (Team et al., 2025), Llama-3.2-Vision-11B (Meta, 2024), MiniCPM-o 2.6 (Yao et al., 2024), InternVL2.5 series (Wang et al., 2024b; Chen et al., 2024c), LLaVA-v1.6-vicuna-7b (Liu et al., 2024), Qwen2.5-VL series (Bai et al., 2025). Details of these models are listed in Table 11 in Appendix D.
@@ 4.2 Main Results
Overall Performance. Across all evaluated models, InternVL2.5-8B-MPO-AWQ achieves the best overall accuracy (Accq = 63.32%), followed closely by InternVL2.5-8B-MPO (61.68%) and MiniCPM-o 2.6 (59.45%). Among proprietary models, GPT-4o outperforms the others, obtaining an overall accuracy of 60.13%, slightly higher than GPT-4V (52.70%). In contrast, Claude-3.5-Sonnet and Kimi-VL underperform, showing limited ability in disambiguation tasks.
Results with Different Model Size. Open-source models with scales larger than 7B generally outperform smaller ones, with all top-performing models falling within this range of scale, which is likely to benefit from richer training data and more advanced architectures. In comparison, models smaller than or equal to 7B show a clear performance gap. Although certain models, such as LaVA-Onevision (57.20%) and mPLUG-Owl2-2B (59.29%), perform competitively, most smaller models struggle with complex ambiguities, particularly in semantic and cultural contexts.
Results on Different Categories. We further break down the results by disambiguation categories, and find that InternVL2.5-8B-MPO-AWQ consistently leads in most categories, especially in Homonymy (70.18%), Grammar (64.71%), and Semantics (69.05%). Notably, MiniCPM-o 2.6 excels in the Specialized category (69.47%), suggesting domain knowledge plays a key role. In the Cultural category, which requires understanding cross-cultural references, models like GPT-4o (47.06%) and InternVL2.5-8B-MPO-AWQ (60.00%) show relatively stronger performance. On the other hand, most models perform poorly in the Polysemy and Dual-Ambiguity categories, reflecting the inherent challenges in resolving subtle or cross-modal ambiguities.
@@
5 Analysis and Discussion
5.1 Ablation Study
@@ 5.2 Discussion
5.3 Further Exploration: An Agent-Based Framework for Ambiguity Resolution
Figure 5: Pipeline of our proposed framework. We first identify ambiguity and its significance from the task description, question, and context/image. Then we generate contextual evidence to resolve the ambiguity. Finally, we bridge the context and image through reasoning to produce the final answer.
@@ 5.3 Further Exploration: An Agent-Based Framework for Ambiguity Resolution
Formally, given a task description $\mathcal{D}$, a question $\mathcal{Q}$, and a multimodal ambiguous context $\mathcal{X}$ (e.g., an image and text), the agent-based model $\mathcal{M}$ solves the task through a three-step process as illustrated in Figure 5:
@@ 6 Conclusion
Ethics Statement
We ensured that all images in our dataset comply with ethical and legal standards. A small portion of images were manually synthesized to cover rare scenarios, and 0.5% were personally collected; both underwent strict quality checks to minimize bias. The remainder were obtained from platforms with free usage rights (e.g., Unsplash, Pexels, Pixabay). All images were carefully reviewed to ensure quality, fairness, and copyright compliance. Given their high quality and small proportion, synthetic images are unlikely to affect evaluation outcomes.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62306161, 62276152, 62236011).
Appendix A Dataset Details
A.1 Dataset Distribution
@@
Prompt for GPT-4o evaluation.
I’ll give you an image. Please answer my question based on the image.
Directly select the correct option (A, B, C, D, or E).
Use the following format to answer:
Answer: [ONLY the option letter; not a complete sentence]
Only give me the reply according to this format, don’t give me any other words.
Now, please answer this question.
Question: [QUESTION HERE]
Options: [OPTIONS HERE]
Table 6: Prompt for GPT-4o evaluation.
Model Consistency Score
GPT-4o 96.25
Gemini-2.0-flash 96.25
Qwen2.5-VL-7B-Instruct 97.35
InternVL2.5-8B-MPO-AWQ 97.57
mPLUG-Owl3-7B-240728 95.26
Table 7: Consistency results of different models.
Name Gender Fluent Languages Country/Region Professional Background
author1 Male Chinese, English China Computer Science, Telecommunications, Systems Science
author2 Female English, Chinese England Computer Science, Telecommunications, Finance, Law
author3 Female Malay, English, Chinese Malaysia Computer Science
author4 Male Chinese, English China Computer Science
author5 Female Chinese, English China Computer Science, Economics
author6 Female Chinese, English China Computer Science, Economics, Statistics, Data Science, Humanities and Social Sciences
author7 Male English, Chinese, Japanese China Computer Science, Data Science
author8 Male Chinese, English China Computer Science, Statistics
Table 8: Information about Annotators
Human 75.00 86.59 73.50 86.00 70.09 85.05 78.57 89.29 79.55 89.36 68.29 83.53 81.43 87.12 74.66 86.42
Table 9: Human Evaluation Results. For humans, the two experimental settings show little difference, so we chose the more challenging with confusing options setting for human evaluation.
A.2 Information about Annotators
In our study, the annotators are the co-authors of the paper. We opted for this arrangement to ensure careful and consistent evaluation, given the nuanced nature of the tasks involving ambiguity and cross-linguistic interpretation. Below, we provide detailed information about the annotators, including their linguistic fluency, country/region of residence, professional background, and gender. Table 8 shows the information about annotators. It is important to note that all the images used in the dataset were carefully selected and manually reviewed. The selection process involved 8 data curators, who cross-checked the images to ensure quality and relevance. Only those images that received unanimous approval from all curators were included in the dataset. This approach ensured that the images met the necessary standards for consistency and quality.
A.3 Potential Risks Analysis
As for the details of data collection, we estimate that approximately 2.3% of the images are AI-generated and 0.5% of the images are taken personally, while the rest are crawled from online sources (e.g., Unsplash, Pexels, Pixabay).
Regarding the potential impact of synthetic images on the evaluation of multimodal ambiguity resolution, all synthetic images were carefully crafted to closely resemble real-world scenes and passed strict manual quality checks. They were designed to supplement rare or hard-to-collect cases, enhancing benchmark coverage without introducing bias. Given their high quality and small proportion in the dataset, any potential impact on model performance evaluation is expected to be negligible.
Appendix B Human Evaluation
B.1 Results of Human Evaluation
We have conducted a more comprehensive human evaluation. Specifically, we conducted evaluations with 3 human evaluators for Chinese and English, and 2 human evaluators for Malay. The final evaluation scores were averaged across the evaluators for each language. These new results provided a concrete basis for comparison against the current MLLMs performance. Table 9 is the full breakdown of human performance across various categories. For humans, the two experimental settings show little difference, so we chose the more challenging with confusing options setting for human evaluation. As shown in the table, our benchmark features higher difficulty and a larger dataset. The evaluation scores of human evaluators are significantly higher than those of large models, which highlights a strong gap between human-level performance and current MLLMs. Specifically, the best overall $\mathrm{Acc_{p}}$ of current MLLMs is only 37.32, but human $\mathrm{Acc_{p}}$ reaches 74.66.
B.2 Consistency between GPT and Human Evaluation
Due to the diverse and often unpredictable nature of large model outputs, it is difficult to strictly evaluate them based on exact match criteria, as this approach would not always be accurate. Additionally, human evaluation is highly time-consuming and labor-intensive, especially when dealing with large datasets. So, we adopt GPT-based evaluation when assessing model performance.
To ensure the reliability of GPT-based evaluation, we conducted a consistency analysis between human and GPT judgments across multiple models by using prompt in Table 6. The results in Table 7 show high agreement rates, where the GPT-based binary judgments aligned with human evaluation in over 94% of the cases, supporting the feasibility of using GPT-based evaluation in place of manual evaluation.
Appendix C Prompt for Model Inference
@@
Table 10 presents the detailed prompts used during model testing. In the main experiments, we employed English prompts; additionally, we conducted ablation studies using Chinese and Malay prompts to evaluate the impact of different evaluation languages on the experimental results. The table also lists the prompts used in the three ablation settings: $\mathcal{(Q,I)}$, $\mathcal{(C,Q)}$, and $\mathcal{(C)}$, where $\mathcal{(Q,I)}$ uses the Question and Image as input, $\mathcal{(C,Q)}$ uses the Context and Question, and $\mathcal{(C)}$ uses only the Context.
@@
Appendix D Evaluated Models
@@
Deepseek-VL-Tiny (Lu et al., 2024) $7\mathrm{B}$ SigLIP DeepSeek LLM MLP
@@
Qwen2.5-VL-3B-Instruct (Bai et al., 2025) ViT Qwen2.5 LLM MLP
Qwen2.5-VL-7B-Instruct (Bai et al., 2025) ViT Qwen2.5 LLM MLP
@@
We evaluate a total of 19 models covering a range of scales and architectures. Our evaluated proprietary models include GPT-4V (OpenAI, 2023), GPT-4o (OpenAI, 2024), Gemini (Gemini Team et al., 2023), and Claude-3.5-Sonnet (Anthropic, 2024). For open-source models, we include Deepseek-VL-Tiny (Lu et al., 2024), Kimi-VL (Team et al., 2025), Llama-3.2-Vision-11B (Meta, 2024), MiniCPM-o 2.6 (Yao et al., 2024), InternVL2-8B (Chen et al., 2024d), InternVL2.5-8B-MPO (Wang et al., 2024b), InternVL2.5-8B-MPO-AWQ (Chen et al., 2024c), LLaVA-v1.6-vicuna-7b (Liu et al., 2024), Qwen2.5-VL-3B-Instruct (Bai et al., 2025), Qwen2.5-VL-7B-Instruct (Bai et al., 2025). Details of these models are listed in Table 11. Table 11 presents a comprehensive overview of the Multimodal Large Language Models (MLLMs) evaluated in our benchmark. The models are categorized into two groups: API-based models and open-source models. For each model, we list its parameter size category (greater than or less than 7 billion), the vision encoder architecture, the underlying language model (LLM) backbone, and the employed vision-to-language (V2L) adapter. API-based models such as GPT-4V and Gemini do not publicly disclose architectural details, while open-source models span a variety of encoders (e.g., SigLIP, CLIP ViT, InternViT), LLM backbones (e.g., Llama, Qwen, InternLM), and adapter types (e.g., MLP, Linear, XAttn LLM). This table highlights the diversity in architectural design choices across MLLMs.
Table 12 reports results for Chinese and Malay prompts under two settings: with and without confusing options. The four sections present detailed model performance for each language and setting combination.
@@
In Chinese, with confusing options
@@
In Malay, without confusing options
@@
In Malay, with confusing options
@@
Table 12: Ablation study in four settings: Chinese and Malay prompts, each with and without confusing options.
As shown in Table 12, models achieve the highest performance without confusing options, with InternVL2.5-8B-MPO-AWQ reaching 65.82 for Chinese prompts and 65.27 for Malay prompts. When confusing options are introduced, overall accuracy drops noticeably: for Chinese prompts, the top score decreases by 7.72 (65.82 → 58.10), while for Malay prompts, the drop is even larger at 11.69 (65.27 → 53.58). This indicates that confusing options substantially increase task difficulty. Across both languages, Chinese prompts perform slightly better than Malay prompts, though the gap remains small (0.55 without confusing options and 4.41 with confusing options). InternVL2.5-8B-MPO-AWQ consistently achieves the best results across all settings.
Appendix E More Cases
We present a comprehensive set of additional cases from the MUCAR dataset to further illustrate the performance of Multimodal Large Language Models (MLLMs). Specifically, Figure 6, Figure 7, Figure 8, Figure 9, Figure 10, Figure 11, Figure 12, Figure 13 display model outputs across various scenarios. For clarity and balance, we select five representative cases from each of the seven predefined categories. In these visualizations, incorrect responses generated by the models are clearly marked in red to allow easy identification of errors.
To support a deeper understanding of the visual and contextual challenges within each case, we also include detailed explanations that highlight the key ambiguities present in the images. These annotations are intended to help readers recognize why a particular question might be difficult to answer correctly, either for a model or a human.
However, it is crucial to emphasize that these explanatory notes were not accessible to either the MLLMs or the human participants during the question-answering process. Both models and human volunteers provided their responses without the benefit of additional contextual guidance, ensuring a fair and unbiased assessment of performance. This setup allows us to isolate and better evaluate the inherent reasoning and perception capabilities of the models in comparison to human interpretation.
Overall, these additional cases offer further insight into the specific limitations and strengths of MLLMs when dealing with multimodal ambiguity, reinforcing the broader findings of our evaluation.